solve by elimination
5x + 5y = -13
7x - 2y = 17

Found 2 solutions by Tobiasz, Alan3354:
Answer by Tobiasz(54)

ok: if any questions about the problem once i solve it email me at tobiasz.chris@gmail.com!!
-7(5x + 5y = -13)
5(7x - 2y = 17)
-35x - 35y = 91
35 + 10y = 85
______________________
45y = 176
divide both of 45
y+=+3.91}
5x + 5y = -13
5x + 5(3.91) = -13
5x + 19.55 = -13
5x = -32.55
divide both of 5
x+=+-6.51
5x +5y = -13
5 ( 6.51) + 5 ( 3.91) = -13
-32.55 + 19.55 = -13
-13+=+-13
OK




Answer by Alan3354(69443)

solve by elimination
5x + 5y = -13
7x - 2y = 17
-------------
10x + 10y = -26 1st eqn x 2
35x - 10y = 85 2nd eqn x 5
--------------- Add
45x = 59
x = 59/45
----------
5x + 5y = -13
59/9 + 5y = -13 = -117/9
5y = (-117 - 59)/9 = -176/9
y = -176/45